# RFWave: Multi-band Rectified Flow for Audio Waveform Reconstruction

### v2

@@ 5 Results
Table 4: Model footprint and synthesis speed. xRT stands for the speed at which the model can generate speech in comparison to real-time. A higher xRT value signifies that the model is capable of producing speech quicker than real-time, with a value of 1.0 representing the speed of real-time.
Model Parameters GPU xRT($\uparrow$) Sampling steps
RFWave 18.1 M 97.01 10
Vocos(ISTFT) 13.5 M 2078.20 1
PriorGrad 2.6 M 16.67 6
FreGrad 1.7 M 7.50 50
RFWave(CFG2) 18.1 M 64.87 10
MBD 411.0 M 4.82 20
6 Conclusion and Discussion
@@
In this study, we propose RFWave, a multi-band Rectified Flow approach for audio waveform reconstruction. The model has been carefully designed to overcome the latency issues associated with diffusion models. RFWave stands out for its ability to generate complex spectrograms by operating at the frame level, processing all subbands concurrently. This concurrent processing significantly enhances the efficiency of the waveform reconstruction process. The empirical evaluations conducted in this research have demonstrated that RFWave achieves exceptional reconstruction quality. Moreover, it has shown superior computational efficiency by generating audio at a speed that is 97 times faster than real-time.
It would be relatively easy to implement the widely-used cascade pipeline to develop a Text-to-Speech (TTS) system. This involves mapping text features to Mel-spectrograms and then Mel-spectrograms to complex spectrograms using Rectified Flow for both stages. Nevertheless, it is more advantageous to map text features directly to complex spectrograms, especially in the context of rapidly evolving large-scale TTS models. Large-scale TTS models typically incorporate extensive corpora, and eliminating one stage of processing can significantly reduce computational resource requirements. Moreover, this direct approach minimizes the potential inconsistencies that may occur between the two stages. The infilling capabilities of Rectified Flow enable it to handle diverse functions similar to those undertaken by prominent large-scale TTS models, for example, replicating the speaker’s voice and speech style from a provided audio prompt [42, 43]. We believe this approach warrants further investigation and plan to explore it in future work.
Ethical Concerns
Our method, though not categorized as generative AI, is designed to be compatible with advanced techniques such as those in [17], improving the realism of synthesized speech. However, this also poses risks like the potential for creating convincing deepfakes and voice phishing scams.
@@ A.1 Implementation Details
The RFWave backbone contains 8 ConvNeXtV2 blocks. Within each ConvNeXtV2 block, the depth-wise convolutional layer featuring a large kernel utilizes a kernel size of 7 and has a channel dimension of 512. The first and last 1x1 point-wise convolutional layers in the sequence possess channel dimensions of 512 and 1536, respectively.
@@ A.2 Computational Resource Required for the Experiments
A.3 Spectrogram Examples
@@
Figure A.3 exhibits spectrogram examples generated by different models utilizing the LibriTTS dataset. RFWave and Vocos both generate clear high-frequency harmonics, while PriorGrad and FreGrad result in blurred high-frequency harmonics.
Figure A.4 displays spectrogram examples generated by diverse models from the Jamendo dataset. Both RFWave and PriorGrad generate commendable spectrograms, with RFWave edging out slightly in the high-frequency range. Interesting artifacts are observed in the spectrograms produced by Vocos and FreGrad. Vocos tends to generate very sharp and straight harmonics, leading to metallic sound. In contrast, FreGrad produces spectrograms with vertical lines, resulting in a sound that is jittery and discontinuous.
@@
A.4 Tables
@@
Speech Expresso-test 50 From the Expresso dataset test set [44], randomly select 50 samples, segmenting any audio longer than 30 seconds into 10-second clips beforehand.
HiFiTTS-test 50 Randomly select 50 samples from the test set of the Hi-Fi TTS dataset [45].
LibriTTS-test 50 Randomly select 50 samples from the test set of LibriTTS [31].
Aishell3-test 50 Randomly select 50 samples from the test set of the Aishell3 dataset [46].
JVS 50 Randomly select 50 samples from the entire JVS Corpus. [47].
CML-TTS-test 50 Randomly select 50 samples from the test set of the CML-TTS dataset [48].
Vocal Musdb-test-vocal 60 Segment the vocal track audio from the Musdb test set [49] into 10-second clips, then randomly select 60 samples.
CSD 20 Segment all audio in the CSD [50] into 10-second clips, then randomly select 20 samples.
Opencpop 20 Randomly select 20 samples from the Opencpop dataset [29].
ChineseOpera-monophonic 60 From the monophonic subset of the Chinese Opera Singing Dataset [51], segment any audio longer than 30 seconds into 10-second clips, then randomly select 60 samples.
JVS-Music 60 Randomly select 60 samples from the JVS-Music Corpus [52].
RAVDESS 60 Randomly select 60 samples from the song subset of the RAVDESS corpus [53].
Ccmusic-demo-vocal 20 From the demo audios of the Ccmusic Dataset [54], select the vocal parts. Then, randomly select 20 samples from these for the test set, segmenting any audio longer than 30 seconds into 10-second clips beforehand.
Sound Effect ESC-50 150 Randomly select 150 samples from the ESC-50 [55].
Musdb-test-accompaniment 40 Segment the accompaniment track audio from the test set of the Musdb dataset [49] into 10-second clips, then randomly select 40 samples from these clips.
Musdb-test-mixture 20 Segment the mixture track audio from the test set of the Musdb dataset [49] into 10-second clips, then randomly select 20 samples from these clips.
ChineseOpera-polyphonic 20 From the polyphonic subset of the Chinese Opera Singing Dataset [51], segment any audio longer than 30 seconds into 10-second clips, then randomly select 20 samples.
OpenMIC-test 40 Randomly select 40 samples from the test set of OpenMix 2018 [56].
Ccmusic-demo-music 30 From the demo audios of the Ccmusic Dataset [54], select the non-vocal parts. Then, randomly select 20 samples from these for the test set, segmenting any audio longer than 30 seconds into 10-second clips beforehand.
Table A.5: MOS and Objective evaluation metrics for RFWave, Vocos, PriorGrad and FreGrad across various datasets.

### v3

@@ 4.5 Inference Speed
Table 7: Model footprint and synthesis speed. xRT stands for the speed at which the model can generate speech in comparison to real-time.
Model Parameters (M) GPU xRT $\uparrow$ GPU Memory (MB) Sampling steps
RFWave 18.1 162.59 780 10
BigVGAN 107.7 72.68 1436 1
Vocos(ISTFT) 13.5 2078.20 590 1
PriorGrad 2.6 16.67 4976 6
FreGrad 1.7 7.50 2720 50
MBD 411.0 4.82 5480 20
RFWave(44.1KHz) 20.5 152.58 902 10
BigVGAN(44.1KHz) 116.5 39.03 1740 1
5 Conclusion
@@
In this study, we propose RFWave, a multi-band Rectified Flow approach for audio waveform reconstruction. The model has been carefully designed to overcome the latency issues associated with diffusion models. RFWave stands out for its ability to generate complex spectrograms by operating at the frame level, processing all subbands concurrently. This concurrent processing significantly enhances the efficiency of the waveform reconstruction process. The empirical evaluations conducted in this research have demonstrated that RFWave achieves exceptional reconstruction quality. Moreover, it has shown superior computational efficiency by generating audio at a speed that is 160 times faster than real-time, comparable to GAN-based methods, making it practical for real-world applications.
@@ A.1 Implementation Details
The RFWave backbone contains 8 ConvNeXtV2 blocks. Within each ConvNeXtV2 block, the depth-wise convolutional layer featuring a large kernel utilizes a kernel size of 7 and has a channel dimension of 512. The first and last 1x1 point-wise convolutional layers in the sequence possess channel dimensions of 512 and 1536, respectively. In Subsection 4.4, the half-size variant contains 8 ConvNeXtV2 blocks with point-wise convolutional layers of 384 and 1152, while the double-size variant features 16 ConvNeXtV2 blocks with point-wise convolutional layers of 512 and 1536.
@@ A.2 Computational Resource Required for the Experiments
A.3 Deviding into Subbands
The complex spectrogram is circularly padded in the feature dimension with a size of $(d_{ol},d_{ol}-1)$, where $d_{ol}$ signifies the overlap size. The dimension of each subband’s main section, denoted as $d_{m}$, is calculated by dividing $d-1$ by the total number of subbands. Here, $d$ represents the dimension of the complex spectrograms. The last subband is an exception, having an extra feature dimension but one less padding dimension. Subbands are extracted by applying a sliding window along the feature dimension, where the window has a size of $d_{m}+2d_{ol}$ and shifts by $d_{m}$ [^4] The feature dimension of each subband, represented as $d_{s}$, equals $2(d_{m}+2d_{ol})$, accounting for the interleave of real and imaginary parts..
A.4 STFT Loss
Spectral convergence (SC) loss
${\||X_{1}|-|\overset{\sim}{X_{1}}|\|_{F}}/{\||X_{1}|\|_{F}},$ (8)
where $X_{1}$ is the ground truth and $\overset{\sim}{X_{1}}$ is the estimated complex spectrogram using (7). The Frobenius norm, $\|\cdot\|_{F}$, applied over time and frequency, emphasizes large spectral components in SC loss.
Log-scale STFT-magnitude loss
$\|\log(|X_{1}|+\epsilon)-\log(|\overset{\sim}{X_{1}}|+\epsilon)\|_{1},$ (9)
where the $L^{1}$ norm, $\|\cdot\|_{1}$, along with a small constant $\epsilon$, is used to accurately capture small-amplitude components in the log-scale STFT-magnitude loss.
A.5 Selecting Time Points for Euler Method
Straightness is calculated from a single batch with a size of 96. The Euler method uses 100 equal interval steps to estimate the integral in the straightness definition. Time points are selected such that the increase in straightness remains consistent across each interval. An example with 10 Euler steps is provided in Figure A.2. The time points are calculated only once per model, resulting in negligible computational load.
Figure A.2: Deviation (top) and straightness (bottom) over time: red dots mark Euler method time points, with constant increase in straightness across intervals.
A.6 Spectrogram Examples for Comparison with Diffusion-based Method
@@
Figure A.4 exhibits spectrogram examples generated by different models utilizing the LibriTTS dataset. RFWave generate clear high-frequency harmonics, while PriorGrad and FreGrad result in blurred high-frequency harmonics.
Figure A.5 displays spectrogram examples generated by diverse models from the Jamendo dataset. Both RFWave and PriorGrad generate commendable spectrograms, with RFWave edging out slightly in the high-frequency range.
@@
A.7 Spectrogram Examples for Comparison with GAN-based Method
Figure A.6 illustrates that RFWave, BigVGAN and Vocos are capable of generating high-quality spectrograms when applied to the LibriTTS dataset. This demonstrates the effectiveness of each approach in handling in-distribution data. The spectrograms exhibit well-defined harmonics and minimal artifacts, highlighting the robustness of the models within their familiar domain.
Figure A.7 presents spectrogram examples generated on the MUSDB18 dataset, with the models having been trained on the LibriTTS dataset. A notable observation is that both BigVGAN and Vocos exhibit a tendency to produce horizontal lines in the high-frequency regions of the spectrograms. These artifacts are perceptually problematic, as they often result in a metallic sound quality that detracts from the naturalness of the audio. In stark contrast, RFWave consistently generates clear and well-defined high-frequency harmonics. This capability underscores RFWave’s superior ability to maintain spectral fidelity, even when applied to out-of-distribution data, thereby enhancing the overall audio quality and realism.
(a) Ground truth
(b) RFWave
(c) BigVGAN
(d) Vocos
Figure A.6: Examples of spectrograms from LibriTTS
(a) Ground truth
(b) RFWave
(c) BigVGAN
(d) Vocos
Figure A.7: Examples of spectrograms from MUSDB18
(a) Ground truth
(b) w/ STFT Loss
(c) w/o STFT Loss
Figure A.8: The effect of STFT loss. There are some vertical patterns in the spectrogram of waveforms generated by a model without STFT loss, as highlighted by the green rectangular.
(a) Ground truth
(b) w/ overlap Loss
(c) w/o overlap Loss
Figure A.9: The effect of overlap loss. Omitting it results in a noticeable transition between subbands, as clearly illustrated by the green arrows.
A.8 Tables
@@
Vocos https://github.com/gemelo-ai/vocos
BigVGAN https://github.com/NVIDIA/BigVGAN
@@
Speech Expresso-test 50 From the Expresso dataset test set (Nguyen et al., 2023), randomly select 50 samples, segmenting any audio longer than 30 seconds into 10-second clips beforehand.
HiFiTTS-test 50 Randomly select 50 samples from the test set of the Hi-Fi TTS dataset (Bakhturina et al., 2021).
LibriTTS-test 50 Randomly select 50 samples from the test set of LibriTTS (Zen et al., 2019).
Aishell3-test 50 Randomly select 50 samples from the test set of the Aishell3 dataset (Shi et al., 2020).
JVS 50 Randomly select 50 samples from the entire JVS Corpus. (Takamichi et al., 2019).
CML-TTS-test 50 Randomly select 50 samples from the test set of the CML-TTS dataset (Oliveira et al., 2023).
Vocal Musdb-test-vocal 60 Segment the vocal track audio from the Musdb test set (Rafii et al., 2017) into 10-second clips, then randomly select 60 samples.
CSD 20 Segment all audio in the CSD (Choi et al., 2020) into 10-second clips, then randomly select 20 samples.
Opencpop 20 Randomly select 20 samples from the Opencpop dataset (Wang et al., 2022).
ChineseOpera-monophonic 60 From the monophonic subset of the Chinese Opera Singing Dataset (Black et al., 2014), segment any audio longer than 30 seconds into 10-second clips, then randomly select 60 samples.
JVS-Music 60 Randomly select 60 samples from the JVS-Music Corpus (Tamaru et al., 2020).
RAVDESS 60 Randomly select 60 samples from the song subset of the RAVDESS corpus (Livingstone & Russo, 2018).
Ccmusic-demo-vocal 20 From the demo audios of the Ccmusic Dataset (Liu & Li, 2021), select the vocal parts. Then, randomly select 20 samples from these for the test set, segmenting any audio longer than 30 seconds into 10-second clips beforehand.
Sound Effect ESC-50 150 Randomly select 150 samples from the ESC-50 (Piczak, 2015).
Musdb-test-accompaniment 40 Segment the accompaniment track audio from the test set of the Musdb dataset (Rafii et al., 2017) into 10-second clips, then randomly select 40 samples from these clips.
Musdb-test-mixture 20 Segment the mixture track audio from the test set of the Musdb dataset (Rafii et al., 2017) into 10-second clips, then randomly select 20 samples from these clips.
ChineseOpera-polyphonic 20 From the polyphonic subset of the Chinese Opera Singing Dataset (Black et al., 2014), segment any audio longer than 30 seconds into 10-second clips, then randomly select 20 samples.
OpenMIC-test 40 Randomly select 40 samples from the test set of OpenMix 2018 (Humphrey et al., 2018).
Ccmusic-demo-music 30 From the demo audios of the Ccmusic Dataset (Liu & Li, 2021), select the non-vocal parts. Then, randomly select 20 samples from these for the test set, segmenting any audio longer than 30 seconds into 10-second clips beforehand.
Table A.5: MOS and Objective evaluation metrics for RFWave, PriorGrad and FreGrad across various datasets. [^6] The MOS of the ground truth on LibriTTS differs from that presented in Table 2 because these two MOS tests were conducted separately.